Circular No. 6525

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cc IN NGC 5673 S. Nakano, Sumoto, Japan, reports the discovery by Shunji Sasaki,

Hasaki-machi, Ibaraki-ken, of an apparent supernova (m_{pg} = 15) on two 4-min T-Max 400 films taken on Dec. 19.769 and 19.773 UT with a 0.20-m f/4.8 reflector. Sasaki reports the new star as being ∼ 20′′ northwest of the center of NGC 5673, noting that a brighter foreground star (mag 12.5) is located ∼ 40′′ northwest of the galaxy’s nucleus. Sasaki’s patrol films in 1996 show no star at the location of SN 1996cc. R. Kushida, Yatsugatake South Base Observatory, confirms the presence of SN 1996cc (at mag 14.9) on an unfiltered CCD frame taken on Dec. 20.733; nothing was present at this location on a similar frame taken on 1996 May 25. Y. Kushida provides the following measurement of SN 1996cc from the Dec. 20 exposure: α = 14^{h}31^{m}28^{s}.45, δ = +49^{o}57′43′′.8 (equinox 2000.0), which is 25′′ west and 14′′ north of the galaxy’s center. A star of mag 15 is located at position end figures 30^{s}.0, 35′′.

GRS 1915+105 C. R. Robinson, B. A. Harmon, W. S. Paciesas, K. J. Deal, S. N. Zhang,

M. L. McCollough, and C. A. Wilson, for the Compton Observatory BATSE Team; E. B. Waltman, Naval Research Laboratory; and E. Morgan, Massachusetts Institute of Technology, report: “The galactic superluminal jet source GRS 1915+105 has reached the highest flux levels detected in prior outbursts from BATSE, for data averaged over one day. Previously, three strong outbursts were observed, with peaks averaging around 350 mCrab (20–100 keV) and characterized by a power law with photon index between −3.0 and −2.5. This latest hard x-ray outburst has lasted 8 months, rising slower to its peak flux than the other outbursts and exhibiting, until recently, a spectrum steeper than −3.0. Starting around Nov. 24, the source has remained above 350 mCrab, reaching at times 500 mCrab. The spectrum is now harder, with an index of −2.82 ± 0.04. During this same interval, the RXTE/ASM data (2–10 keV) show average countrates and variability at around their lowest levels thus far detected from this source. Public-domain data from the NSF-NRAO-NASA Green Bank Interferometer, however, reveal the source to be flaring during this interval of high, hard x-ray flux. The largest flare was detected on Dec. 6 at a flux level of nearly 150 mJy (2.25 GHz). Additional observations of this source are encouraged.”

1996 December 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green